# A Particle-in-Cell Simulation Framework for Thomson Scattering Analysis in Inertial Confinement Fusion

### v1

@@ 1 INTRODUCTION
$P_{s}(\vec{R},\omega_{s})d\Omega d\omega_{s}=\frac{P_{i}e_{0}^{2}}{A2\pi}d\Omega d\omega_{s}\left(1+\frac{2\omega}{\omega_{i}}\right)NS(\vec{k},\omega),$ (1)
where $P_{i}$ denotes the power of the incident (probe) beam, $r_{0}$ is the classical electron radius, and $A$ is the cross-sectional area of the scattering volume. $N$ represents the number of electrons within this volume. $\omega_{i}$ and $\omega_{s}$ are the frequencies of the incident and scattered light, respectively. $\omega$ and $\vec{k}$ denote the frequency and wave vector of the electron density fluctuations modes that satisfy the wave–wave matching conditions:
@@
Where $\chi_{e}$ and $\chi_{i}$ are the susceptibilities of electron and ion respectively, $\epsilon=1+\chi_{e}+\chi_{i}$ is the plasma dielectric function, $f_{e}$ and $f_{i}$ represent the electron and ion velocity distribution function respectively. These distributions embody key plasma parameters, including density, temperature, and flow velocity. Consequently, by fitting measured thermal CTS spectra, one can infer these plasma properties[15, 16, 17, 18, 19, 20, 21, 22, 23].
@@
Particle-in-cell (PIC) simulation is a first-principles method that self-consistently solves Maxwell’s equations and tracks particle motion under electromagnetic forces [27], making it well-suited for studying kinetic effects in nonthermal or collisional plasmas. PIC simulations have been used to examine non-collective TS[28]. For CTS, pioneer investigations have proposed 1D and 2D PIC methods to study nonthermal plasmas [29] along certain directions. While recent experiments seeking to obtain detailed plasma properties or to diagnose non-Maxwellian plasmas via collective Thomson scattering (CTS) often rely on collecting scattered light from multiple angles [30, 31], high-resolution, angle- and frequency-resolved data are still required to guide experimental design. To integrate experimental TS measurements, development of numerical simulation methods with sufficient angular and spectral resolution is essential.
@@
The remainder of this paper is organized as follows: Section II describes the numerical method of obtaining CTS signals with high resolution in angular and frequency space; Section III presents the capability of our method, including numerical results of CTS spectra with different level of ion-ion collisional effects; Section IV focuses on SCTS, in which the new phenomenon, drawing attention to a newly observed phenomenon in which scattered light results from perturbations with mismatched wave vectors; Section V summarizes the paper and outlines directions for future work on studying ICF plasmas using TS diagnostics.
2 SIMULATION PROCEDURE TO OBTAIN THOMSON SCATTERING SIGNAL
@@ 2.1 PIC simulation setup
A probe beam is injected from the left boundary and propagates along the x-direction, as illustrated in Fig.1. It has a frequency of $\omega_{i}=\tfrac{4}{3}\omega_{0}$ (corresponding to a 263 nm wavelength) and is linearly polarized perpendicular to the xy-plane. The transverse profile of the beam is Gaussian, with a peak intensity $I=9\times 10^{12}\,\mathrm{W/cm^{2}}$ and a full width at half maximum (FWHM) varying from $100\,c/\omega_{0}$ ($5.6\,\mu m$) to $400\,c/\omega_{0}$ ($22.3\,\mu m$).
The spatial grid resolution is $\Delta x=0.25\,c/\omega_{0}$ ($0.014\,\mu m$), and the temporal step size is $\Delta t=0.17\,\omega_{0}^{-1}$ ($0.003\,\text{fs}$). Each grid cell contains 100 macroparticles representing electrons and another 100 for ions. Particle boundary conditions are thermal, while electromagnetic fields satisfy open boundary conditions.
@@ 2.2 Frequency- and angular-resolved TS spectrum
3 ION-ACOUSTIC FEATURES OF THERMAL THOMSON SCATTERING OBTAINED FROM SIMULATIONS
@@ 3.1 Ion-acoustic features of Thomson scattering
For the plasma conditions described in Section II, the frequency of ion-acoustic peaks $\omega_{s}$ in the TS spectrum can be calculated using:
@@
Following the method mentioned in Section II.C, we obtain low-frequency part of Thomson scattering spectra at different scattering angles of $60^{\circ}\pm 1^{\circ}$, $90^{\circ}\pm 1^{\circ}$, $120^{\circ}\pm 1^{\circ}$, and $150^{\circ}\pm 1^{\circ}$, as shown in Fig.2. To reduce spectral noise, second-order smoothing with a window size of $3.0\times 10^{-4}\omega_{0}$ is applied iteratively to these spectra until their shapes converge. The simulated TS spectra are in good agreement with the theoretical profiles calculated from Eq. (4), which indicates that our method can produce valid TS spectra with high angular and spectral resolution.
@@ 3.2 Achieving higher signal-to-noise ratio
In Sec II.A, we demonstrated that simulations with a domain size of $800\,c/\omega_{0}\times 800\,c/\omega_{0}$ ($44.7\,\mu m\times 44.7\,\mu m$) is adequate to yield ion-acoustic peaks of nearly equal height and a significantly improved signal-to-noise ratio. But the required simulation space may vary when different physical subjects are studied. So a more general approach to improve statistical robustness is required. This is accomplished by conducting multiple simulations under identical macroscopic conditions but with varied microscopic realizations (e.g., different random seeds for initial velocity sampling), and then averaging their resulting spectra.
Figure 4: Simulated Thomson scattering spectrum with different ion-ion collision rate.
@@ 3.3 Capability to simulate ion-ion collision’s influence on Thomson scattering spectrum
Ion–ion collisions modify the ion-acoustic features of the TS spectrum by enhancing collisional damping and reducing Landau damping [24]. Their influence on TS spectrum with wave vector $k$ can be evaluated with the parameter $k\lambda_{ii}$. Here $\lambda_{ii}$ is the ion mean free path due to ion-ion collisions given by $\lambda_{ii}=v_{ti}/\nu_{ii}$, where $v_{ti}$ represents ion thermal velocity and $\nu_{ii}$ denotes ion-ion collision frequency. The condition $k\lambda_{ii}\gg 1$ corresponds to the collisionless cases, $k\lambda_{ii}\sim 1$ to the intermediate cases, and $k\lambda_{ii}\ll 1$ to the strongly collisional cases.
@@ 4.1 Super-thermal Thomson scattering from matched density perturbations with different frequency
The simulations are performed in a domain of $200\,c/\omega_{0}\times 200\,c/\omega_{0}$ ($11.2\,\mu m\times 11.2\,\mu m$), filled with uniform hydrogen plasma. The plasma and probe parameters are the same as those described in Sec.II. As shown in Fig.5(a), two laser beams, each with an intensity of $3.5\times 10^{14}\,\mathrm{W/cm^{2}}$ and a wave vector of $0.93\,\omega_{0}/c$, overlap at an angle of $90^{\circ}$, driving an IAW with a wave vector $k_{\mathrm{iaw}}$ matched to $60^{\circ}$ Thomson scattering. The frequency difference between the two driving beams corresponds to the IAW eigenfrequency, $\omega_{\mathrm{iaw}}=k_{\mathrm{iaw}}c_{s}$. The corresponding SCTS signal at $60^{\circ}$ is shown in Fig. 5(b), where a narrow-band peak appears at the matched frequency $\omega_{i}+\omega_{\mathrm{iaw}}$.
@@ 4.2 Super-thermal Thomson scattering from matched density perturbations with different wave vector
When driving density perturbations through laser beating, the laser overlap region occupies only a small fraction of the total simulation area, leading to inefficient use of computational resources. An alternative method of generating electron density fluctuations is initializing them at the start of the simulation without injecting ions in the simulation domain. Owing to the nature of PIC simulations, which automatically enforce the Poisson equation [27], this initialized electron density distribution is sustained throughout the simulation. This approach can produce stationary perturbations effectively.
@@ 4.3 Thomson scattering from perturbations with a wave vector deviated from matching condition
The SCTS under matching conditions has been presented in the previous sections. When measuring driven plasma waves using SCTS, the scattering geometry should be carefully designed to satisfy the matching condition. However, plasma conditions and the resulting driven plasma waves are often difficult to predict, and the matching conditions may not always be fulfilled in experiments. To investigate a representative scenario relevant to ICF plasmas, where Thomson scattering is employed to diagnose laser-beat–driven ion waves, we perform a PIC simulation to examine this process under a mismatched condition. In the simulation shown in Fig. 7(a), two drive beams with intensity of $10^{14}W/cm^{2}$ and frequency of $\omega_{0}$ is used to generate density perturbations through beating. A probe beam, with a frequency $\omega_{s}=\frac{4}{3}\omega_{0}$ and intensity of $10^{13}W/cm^{2}$, is launched from the lower boundary at $40.5^{\circ}$ clockwise from the y-axis, passing through the region where drive beams overlap. The scattered light is collected along y-axis at scattering angle $\theta_{s}=40.5^{\circ}$. The driven density perturbation has a wave vector $\vec{k}_{beat}$ that slightly differs from the matched wave vector corresponding to the measured scattering angle, $\vec{k}_{match}$. The deviation is quantified as $\vec{k}_{beat}-\vec{k}_{match}=0.08\,\vec{k}_{s}$. The corresponding density perturbation distribution displayed in Fig. 7(b) shows no signal at the matched position. However, in the scattering spectrum shown in Fig. 7(c), a super-thermal Thomson scattering signal with an amplitude significantly higher than thermal TS signal is observed. This result indicates that the observable scattered light reflects not only the amplitude of the matched mode but also contributions from surrounding unmatched modes. This introduces additional challenges in interpreting SCTS signals from driven ion waves.
@@
Figure 7: (a) Distribution of electric field $E_{z}$ in the simulation domain, along with the wave vectors of drive beams ($\vec{k}_{drive,1}$ and $\vec{k}_{drive,2}$), the driven density perturbation ($\vec{k}_{beat}$), the probe ($\vec{k}_{i}$), the collected scatter light ($\vec{k}_{s}$) and the matched density perturbation ($\vec{k}_{match}$). (b) Electron density perturbation in the phase space $(k_{x},k_{y})$. (c) SCTS spectrum along the y-direction ($\theta_{s}=40.5^{\circ}$), the position of signal corresponding to $\vec{k}_{beat}$ is marked by dashed line.
@@
Figure 8: (a) Spatial distribution for the envelope of perturbed current density $\vec{J}_{\rm pert}$, along with the distribution of excited electromagnetic field profiles for (a) $\omega_{\rm pert}=\omega_{0}$, $k_{\rm pert}=\omega_{0}/c$ and (b) $\omega_{\rm pert}=\omega_{0}$, $k_{\rm pert}=0.5\,\omega_{0}/c$.
@@
We provide an explanation for the SCTS emission from unmatched ion density perturbations. We consider the interaction of the probe beam electric field $\vec{E}_{i}\sin(\vec{k}_{i}\cdot\vec{x}-\omega_{i}t)$ with electron density perturbation $\delta n_{e}\sin(\vec{k}_{0}\cdot\vec{x}-\omega_{0}t)$. A current density perturbation can be directly generated via wave beating: $\vec{J}_{pert}=\frac{\vec{E}_{i}e\delta n_{e}}{2m_{e}}(\cos[(\vec{k}_{i}-\vec{k}_{0})\cdot\vec{x}-(\omega_{i}-\omega_{0})t]-\cos[(\vec{k}_{i}+\vec{k}_{0})\cdot\vec{x}-(\omega_{i}+\omega_{0})t])$. This $\vec{J}_{pert}$, with wave vector $\vec{k}_{pert}=\vec{k}_{i}\pm\vec{k}_{0}$ and frequency $\omega_{pert}=\omega_{i}\pm\omega_{0}$, could excite electromagnetic waves[40]. Starting from the Maxwell equation, we obtain the relation between currents and electric fields:
@@
in which $\vec{J}$ includes $\vec{J}_{pert}$, the perturbation current driven by the interaction between the probe beam and the density perturbations, and $\vec{J}_{cond}$, the conduction current arising from the interaction of background electrons with the electric field. For high frequency electromagnetic wave in a non-magnetized plasma, $\vec{J}_{cond}$ satisfies $\frac{\partial\vec{J}_{cond}}{\partial t}=\frac{e^{2}n_{e}}{m_{e}}\vec{E}$. Substitute $\vec{J}=\vec{J}_{pert}+\vec{J}_{cond}$ into Equation (9), we get:
@@
Therefore, the scattered light associated with unmatched ion density perturbations in a plasma can be approximated as the radiation of $\vec{J}_{\mathrm{pert}}$ in vacuum. We insert the following external electron current into a fully vacuum one-dimensional PIC simulation:

### v2

@@ 1 INTRODUCTION
$P_{s}(\vec{R},\omega_{s})d\Omega d\omega_{s}=\frac{P_{i}r_{0}^{2}}{A2\pi}d\Omega d\omega_{s}\left(1+\frac{2\omega}{\omega_{i}}\right)|\hat{s}\times(\hat{s}\times\hat{E}_{io})|^{2}NS(\vec{k},\omega),$ (1)
where $P_{i}$ denotes the power of the incident (probe) beam, $r_{0}$ is the classical electron radius, and $A$ is the cross-sectional area of the scattering volume. The unit vector $\hat{s}$ denotes the direction of the scattered light, and $\hat{E}_{io}$ is the unit polarization vector of the incident probe beam’s electric field. $N$ represents the number of electrons within this volume. $\omega_{i}$ and $\omega_{s}$ are the frequencies of the incident and scattered light, respectively. $\omega$ and $\vec{k}$ denote the frequency and wave vector of the electron density fluctuations modes that satisfy the wave–wave matching conditions:
@@
Where $\chi_{e}$ and $\chi_{i}$ are the susceptibilities of electron and ion respectively, $\epsilon=1+\chi_{e}+\chi_{i}$ is the plasma dielectric function, $f_{e}$ and $f_{i}$ represent the electron and ion velocity distribution function respectively[15]. These distributions embody key plasma parameters, including density, temperature, and flow velocity. Consequently, by fitting measured thermal CTS spectra, one can infer these plasma properties[16, 17, 18, 19, 20, 21, 22, 23, 24].
@@
Particle-in-cell (PIC) simulation is a first-principles method that self-consistently solves Maxwell’s equations and tracks particle motion under electromagnetic forces [28], making it well-suited for studying kinetic effects in nonthermal or collisional plasmas. PIC simulations have been used to examine non-collective TS[30]. Simulations of electromagnetic fluctuations in pair plasmas have demonstrated good agreement with theoretical predictions [29], validating the capability of PIC to reproduce fluctuation physics. For CTS, where electrostatic fluctuations play a dominant role, pioneer investigations have proposed 1D and 2D PIC methods to study nonthermal plasmas [31] along certain directions. While recent experiments seeking to obtain detailed plasma properties or to diagnose non-Maxwellian plasmas via collective Thomson scattering (CTS) often rely on collecting scattered light from multiple angles [32, 33], high-resolution, angle- and frequency-resolved data are still required to guide experimental design. To integrate experimental TS measurements, development of numerical simulation methods with sufficient angular and spectral resolution is essential.
@@
The remainder of this paper is organized as follows: Section 2 describes the numerical procedure to obtain CTS signals with high resolution in angular and frequency space; Section 3 validates our method’s ability to reproduce thermal CTS spectrum from thermal plasma consistent with current theory, including numerical results of CTS spectra with different level of ion-ion collisional effects; Section 4 focuses on SCTS, in which the new phenomenon, drawing attention to a newly observed phenomenon in which scattered light results from perturbations with mismatched wave vectors; Section 5 summarizes the paper and outlines directions for future work on studying ICF plasmas using TS diagnostics.
2 SIMULATION APPROACH TO OBTAIN THOMSON SCATTERING SIGNAL
@@ 2.1 PIC simulation setup
A probe beam is injected from the left boundary and propagates along the x-direction, as illustrated in Fig.1. It has a frequency of $\omega_{i}=\tfrac{4}{3}\omega_{0}$ (corresponding to a 263 nm wavelength) and is linearly polarized perpendicular to the xy-plane to maximize the CTS signal. The transverse profile of the beam is Gaussian, with a peak intensity $I=9\times 10^{12}\,\mathrm{W/cm^{2}}$ and a full width at half maximum (FWHM) varying from $100\,c/\omega_{0}$ ($5.6\,\mu m$) to $400\,c/\omega_{0}$ ($22.3\,\mu m$).
The spatial grid resolution is $\Delta x=0.25\,c/\omega_{0}$ ($0.014\,\mu m$), and the temporal step size is $\Delta t=0.17\,\omega_{0}^{-1}$ ($0.03\,\text{fs}$). Each grid cell contains 100 macroparticles representing electrons and another 100 for ions. Particle boundary conditions are thermal, while electromagnetic fields satisfy VPML boundary conditions[38] which limits EM wave reflectivity at the order of $\sim 10^{-4}$ in our setup.
@@ 2.2 Frequency- and angular-resolved TS spectrum
3 VALIDATION: ION-ACOUSTIC FEATURES OF THERMAL THOMSON SCATTERING OBTAINED FROM SIMULATIONS
@@ 3.1 Ion-acoustic features of Thomson scattering
For the plasma conditions described in Section 2.1, the frequency of ion-acoustic peaks $\omega_{s}$ in the TS spectrum can be calculated using:
@@
Following the plasma conditions in Section 2.1 and time Fourier analysis method described in Section 2.2, we obtain low-frequency part of Thomson scattering spectra at different scattering angles of $60^{\circ}\pm 1^{\circ}$, $90^{\circ}\pm 1^{\circ}$, $120^{\circ}\pm 1^{\circ}$, and $150^{\circ}\pm 1^{\circ}$ from a simulation domain with a size of $800\,c/\omega_{0}\times 800\,c/\omega_{0}$ ($44.7\,\mu m\times 44.7\,\mu m$), as shown in Fig.2. To reduce spectral noise, second-order smoothing with a window size of $3.0\times 10^{-4}\omega_{0}$ is applied iteratively to these spectra until their shapes converge. The simulated TS spectra are in good agreement with the theoretical profiles calculated from Eq. (4), which indicates that our method can produce valid TS spectra with high angular and spectral resolution.
@@ 3.2 Achieving higher signal-to-noise ratio
In the spectra of Fig. 2, we demonstrated that simulations with a domain size of $800\,c/\omega_{0}\times 800\,c/\omega_{0}$ ($44.7\,\mu m\times 44.7\,\mu m$) is adequate to yield ion-acoustic peaks of nearly equal height and a significantly improved signal-to-noise ratio. But the required simulation space may vary when different physical subjects are studied. So a more general approach to improve statistical robustness is required. This is accomplished by conducting multiple simulations under identical macroscopic conditions but with varied microscopic realizations (e.g., different random seeds for initial velocity sampling), and then averaging their resulting spectra.
@@
In addition to ensemble averaging over random seeds, increasing the number of particles per cell (PPC) is a common technique for improving the signal-to-noise ratio in PIC simulations[41, 42, 43]. While we have employed this method, we find it to be relatively inefficient: raising the PPC reduces both the signal and the statistical noise by similar factors, leaving the overall signal-to-noise ratio largely unchanged. Our tests show that varying the PPC by a factor of 16 results in minimal signal-to-noise ratio improvement. Further details are provided in Appendix A. The insensitivity of signal-to-noise ratio of Thomson scattering signal to PPC could be attribute to the fact that Thomson scattering is related to the ensemble average of density perturbations, while a simulation with an increased PPC does not necessarily enhance the sampling of the fluctuation modes, and therefore contribute little to the ensemble average.
@@ 3.3 Capability to simulate ion-ion collision’s influence on Thomson scattering spectrum
Ion–ion collisions can modify the ion-acoustic features of the TS spectrum by enhancing collisional damping and reducing Landau damping [25]. Their influence on TS spectrum with wave vector $k$ can be evaluated with the parameter $k\lambda_{ii}$. Here $\lambda_{ii}$ is the ion mean free path due to ion-ion collisions given by $\lambda_{ii}=v_{ti}/\nu_{ii}$, where $v_{ti}$ represents ion thermal velocity and $\nu_{ii}$ denotes ion-ion collision frequency. The condition $k\lambda_{ii}\gg 1$ corresponds to the collisionless cases, $k\lambda_{ii}\sim 1$ to the intermediate cases, and $k\lambda_{ii}\ll 1$ to the strongly collisional cases.
Figure 4: Simulated Thomson scattering spectra with different ion-ion collision rate.
@@
In addition, in the above work we intentionally retain only ion-ion collisions to enable a direct and uncluttered comparison with existing theoretical models that specifically account for this process. This approach allows us to cleanly demonstrate the capability of our PIC-based method to reproduce CTS spectra under varying collision strengths. However, the ion-electron and electron-electron collisions should also influence the CTS spectrum. As noted in prior work [46, 47], ion-electron collisions modify the damping rate of ion acoustic waves (IAWs). Similarly, electron-electron collisions can affect the electron contribution to IAW Landau damping. Under conditions where ion-ion collisions are significant, these additional processes may also become non-negligible. While a comprehensive treatment of all collision types is beyond the scope of this initial methodological study, a detailed investigation of multi-species collisional impacts on CTS is planned for future work.
@@ 4.1 Super-thermal Thomson scattering from matched density perturbations with different frequency
The simulations are performed in a domain of $200\,c/\omega_{0}\times 200\,c/\omega_{0}$ ($11.2\,\mu m\times 11.2\,\mu m$), filled with uniform hydrogen plasma. The plasma and probe parameters are the same as those described in Section 2.1. As shown in Fig.5(a), two laser beams, each with an intensity of $3.5\times 10^{14}\,\mathrm{W/cm^{2}}$ and a wave vector of $0.93\,\omega_{0}/c$, overlap at an angle of $90^{\circ}$, driving an IAW with a wave vector $k_{\mathrm{iaw}}$ matched to $60^{\circ}$ Thomson scattering. The frequency difference between the two driving beams corresponds to the IAW eigenfrequency, $\omega_{\mathrm{iaw}}=k_{\mathrm{iaw}}c_{s}$. The corresponding SCTS signal at $60^{\circ}$ is shown in Fig. 5(b), where a narrow-band peak appears at the matched frequency $\omega_{i}+\omega_{\mathrm{iaw}}$.
@@ 4.2 Super-thermal Thomson scattering from matched density perturbations with different wave vector
When density perturbations are driven by laser beating, the region of laser overlap constitutes only a small fraction of the total simulation area. This leads to inefficient use of computational resources. Furthermore, to generate multiple perturbation modes, the laser beating method requires a correspondingly greater number of laser beams. This approach often excites more modes than are necessary, introducing unwanted complexity. For typical cases associated with static density perturbations, an alternative method is to initialize the electrons with a perturbed density distribution $n_{e}(x,y)|_{t=0}$ without any ions injected into the simulation domain. Owing to the nature of PIC simulation [28], it automatically enforces the Poisson equation at any time $t$:
$\nabla\cdot\vec{E}-4\pi\rho=C,$ (7)
where $C$ is a constant determined by initial conditions. When the simulation is initialized with $E|_{t=0}=0$ and $\rho|_{t=0}=-en_{e}(x,y)|_{t=0}$, we obtain $C=4\pi n_{e}(x,y)|_{t=0}$. This form implies that, at any time $t$, the electric potential corresponds to that generated by the instantaneous electron density superimposed on a stationary background ion density equal to $n_{e}(x,y)|_{t=0}$.
@@
Moreover, the angular intensity distributions of the scattered light shown in Fig. 6(b) have finite widths for each scattering angles, with the widths increasing with the scattered angles. This reveals the presence of scattered signals at angles slightly deviating from the exact matching condition. Such scattering under wave-vector-mismatched conditions can be partially attributed to the finite wave-vector broadening of both the probe beam and the driven density perturbation, which relaxes the strict wave-vector-matching requirement by allowing a finite range of wave-vector components to overlap with the matching condition. However, the angular distribution becomes increasingly broader at larger scattering angles, suggesting that wave-vector broadening alone cannot fully account for the observed signal, as analyzed in Appendix B. This behavior therefore points to the presence of additional scattering mechanisms operating under wave-vector-mismatched conditions, which are discussed in Section 4.3.
@@ 4.3 Thomson scattering from perturbations with a wave vector deviated from matching condition
The SCTS under matching conditions has been presented in the previous sections. When measuring driven plasma waves using SCTS, the scattering geometry should be carefully designed to satisfy the matching condition. However, plasma conditions and the resulting driven plasma waves are often difficult to predict, and the matching conditions may not always be fulfilled in experiments.
To investigate a representative scenario relevant to ICF plasmas, where Thomson scattering is employed to diagnose laser-beat–driven ion waves, we perform a PIC simulation to examine this process under a mismatched condition. The simulation is performed in two dimensions (the $xy$-plane) on a domain measuring $400\,c/\omega_{0}\times 400\,c/\omega_{0}$ ($22.4\,\mu m\times 22.4\,\mu m$). The domain is uniformly filled with a CH plasma at an electron density of $n_{e}=0.056\,n_{c}$. Initial electron and ion temperatures are set to $T_{e}=2\,\text{keV}$ and $T_{i}=1\,\text{keV}$, respectively. The spatial resolution is $\Delta x=0.25\,c/\omega_{0}$ ($0.014\,\mu\text{m}$), and the temporal step is $\Delta t=0.17\,\omega_{0}^{-1}$ ($0.03\,\text{fs}$). Each grid cell is populated with 100 macroparticles for electrons and 50 for each ion species. Particle boundary conditions are thermal, while electromagnetic fields satisfy VPML boundary conditions[38].
The simulation employs two drive beams, each with frequency $\omega_{0}$, which beat against each other as illustrated in Fig. 7(a). For the plasma with $n_{e}=0.056\,n_{c}$, the electron plasma frequency is $\omega_{pe}=0.23\omega_{0}$. The electromagnetic dispersion relation yields a wave number $k=0.97\,\omega_{0}/c$ for frequency $\omega_{0}$. As shown in the geometry of Fig. 7(a), drive beam 1 has a wave vector $\vec{k}_{\text{drive},1}\approx(0.04,-0.97)\omega_{0}/c$, oriented $2.5^{\circ}$ counter-clockwise from the negative $y$-axis. Drive beam 2 has $\vec{k}_{\text{drive},2}\approx(-0.81,-0.54)\omega_{0}/c$, oriented $33.5^{\circ}$ counter-clockwise from the negative $x$-axis. The wave vector of the resulting laser-beating density perturbation is $\vec{k}_{\text{beat}}\approx(-0.85,0.43)\omega_{0}/c$.
A probe beam with frequency $\omega_{i}=4\omega_{0}/3$ is launched at an angle of $40.5^{\circ}$ clockwise from the positive $y$-axis. The polarization of the probe beam is set perpendicular to the simulation plane to maximize the Thomson scattering signal. The matched wave vector $\vec{k}_{\text{match}}$ is derived from the frequency-matching condition $\omega_{s}=\omega_{i}+\omega_{\text{beat}}=4\omega_{0}/3$, where the beat frequency $\omega_{\text{beat}}=0$. For the probe frequency of $4\omega_{0}/3$, the magnitude of the wave vector is $|\vec{k}_{i}|\approx 1.31\,\omega_{0}/c$. Specifically, the probe’s wave vector is $\vec{k}_{i}\approx(0.85,0.99)\omega_{0}/c$. Light scattered along the $y$-axis has a wave vector $\vec{k}_{s}\approx(0,1.31)\omega_{0}/c$. The matched wave vector is therefore $\vec{k}_{\text{match}}=\vec{k}_{s}-\vec{k}_{i}\approx(-0.85,0.32)~\omega_{0}/c$. The resulting wave-vector mismatch is $\Delta\vec{k}=\vec{k}_{\text{beat}}-\vec{k}_{\text{match}}\approx(0,0.11)\omega_{0}/c$, which corresponds to $0.08\,|\vec{k}_{s}|$.
The corresponding density perturbation distribution displayed in Fig. 7(b) shows no signal at the matched position. However, in the scattering spectrum shown in Fig. 7(c), a SCTS signal with an amplitude significantly higher than thermal TS signal is observed. The wave vector broadening at finite beam waist is analyzed in Appendix C, and this broadening does not fully account for the observed signal. This result indicates that the observable scattered light reflects not only the amplitude of the matched mode but also contributions from surrounding unmatched modes. This introduces additional challenges in interpreting SCTS signals from driven ion waves.
@@
Figure 7: (a) Distribution of electric field $E_{z}$ in the simulation domain, along with the wave vectors of drive beams $\vec{k}_{drive,1}$ (orange) and $\vec{k}_{drive,2}$ (dark blue)], the driven density perturbation $\vec{k}_{beat}$ (purple)], the probe beam $\vec{k}_{i}$ (blue)], the scattered light $\vec{k}_{s}$ (black)] and the matched density perturbation $\vec{k}_{match}$ (dashed red)]. (b) The spatial spectrum of the electron density perturbation in the phase space $(k_{x},k_{y})$. (c) The SCTS spectrum along $\vec{k}_{s}$ with the peak position marked by the vertical dashed line.
@@
Figure 8: Spatial envelopes of the perturbed current density $\vec{J}_{\rm pert}$ (red), along with the envelopes of excited electromagnetic field (blue) for (a) $\omega_{\rm pert}=\omega_{0}$, $k_{\rm pert}=0.5\omega_{0}/c$, (b) $\omega_{\rm pert}=\omega_{0}$, $k_{\rm pert}=0.8\omega_{0}/c$, (c) $\omega_{\rm pert}=\omega_{0}$, $k_{\rm pert}=1.5\omega_{0}/c$, (d) $\omega_{\rm pert}=\omega_{0}$, $k_{\rm pert}=1.2\omega_{0}/c$ and (e) $\omega_{\rm pert}=\omega_{0}$, $k_{\rm pert}=\omega_{0}/c$.
The occurrence of SCTS from density perturbation modes that do not satisfy the conventional wave-vector matching conditions [Eq. (2)] appears to contradict the standard theory of Thomson scattering, which requires strict frequency and wave-vector matching between the probe beam, the scattered light and the density perturbation. To resolve this puzzle, we must return to the fundamental principle of electromagnetic (EM) wave emission from oscillating electrons. Irrespective of whether the matching conditions are met, the total emitted EM field should always be a coherent superposition of the waves radiated by all oscillating currents induced by the electron motion. Consequently, an analysis of the radiation produced by this current distribution is essential to uncover the underlying physics of SCTS from unmatched modes.
@@
We consider the interaction of the probe beam electric field $\vec{E}_{i}\sin(\vec{k}_{i}\cdot\vec{x}-\omega_{i}t)$ with electron density perturbation $\delta n_{e}\sin(\vec{k}_{0}\cdot\vec{x}-\omega_{0}t)$. A current density perturbation can be directly generated via wave beating:
$\vec{J}_{pert}$ $=\frac{\vec{E}_{i}e\delta n_{e}}{m_{e}}\sin(\vec{k}_{i}\cdot\vec{x}-\omega_{i}t)\sin(\vec{k}_{0}\cdot\vec{x}-\omega_{0}t)$ (8)
$=\frac{\vec{E}_{i}e\delta n_{e}}{2m_{e}}(\cos[(\vec{k}_{i}-\vec{k}_{0})\cdot\vec{x}-(\omega_{i}-\omega_{0})t]-\cos[(\vec{k}_{i}+\vec{k}_{0})\cdot\vec{x}-(\omega_{i}+\omega_{0})t]).$
This $\vec{J}_{pert}$, with wave vector $\vec{k}_{pert}=\vec{k}_{i}\pm\vec{k}_{0}$ and frequency $\omega_{pert}=\omega_{i}\pm\omega_{0}$, could excite EM waves[48]. Starting from the Maxwell equation, we obtain the relation between currents and electric fields:
@@
in which $\vec{J}$ includes $\vec{J}_{pert}$, the perturbation current driven by the interaction between the probe beam and the density perturbations, and $\vec{J}_{cond}$, the conduction current arising from the interaction of background electrons with the electric field. For high frequency electromagnetic wave in a non-magnetized plasma, $\vec{J}_{cond}$ satisfies $\frac{\partial\vec{J}_{cond}}{\partial t}=\frac{e^{2}n_{e}}{m_{e}}\vec{E}$. Substitute $\vec{J}=\vec{J}_{pert}+\vec{J}_{cond}$ into Eq. (9), we get:
@@
Equation (11) describes the growth of an EM wave driven by a perturbative current density, $\vec{J}_{\mathrm{pert}}$, in a plasma with frequency $\omega_{pe}$. If the wave vector $\vec{k}_{\mathrm{pert}}$ and frequency $\omega_{\mathrm{pert}}$ satisfy the EM dispersion relation of the background plasma, the matching condition of Eq. (2) is fulfilled. In this resonant case, the emission is strong because the driven wave corresponds to an eigenmode of the system. Conversely, when the matching condition is not satisfied, the driven wave is not an eigenmode and cannot propagate over long distances in the plasma. Nevertheless, such non-eigenmodes can still give rise to scattered radiation of eigenmodes with finite amplitude. Physically, as these driven non-eigenmodes exit the interaction region between the probe beam and the density perturbations, their oscillating EM fields act as boundary sources that can excite propagating EM eigenmodes at the same frequency in the outer regions.
We demonstrate this mechanism—the excitation of eigenmodes by localized non-eigenmodes—using PIC simulations. It is important to note that, while Eq. (11) describes electromagnetic emission in a plasma, the underlying mechanism is more general and does not inherently depend on the presence of a plasma. To isolate and illustrate the essential physics, we therefore perform the demonstration in vacuum, where no physical particles are present in the simulation domain. In this simplified setup, the perturbative current $\vec{J}_{\mathrm{pert}}$ is imposed directly as an external source. Thus, we insert the following external electron current into a fully vacuum one-dimensional PIC simulation: